A solution of a weak monoprotic base B (0.50 M) and its salt BHCl (0.80 M)...

  1. A solution of a weak monoprotic base B (0.50 M) and its salt BHCl (0.80 M) has a pH = 9.40. What is the value of Kb for this base, B?
Answer #1


use:
pH = 14 - pOH
9.4= 14 - pOH
= 4.6

use:
pOH = pKb + log {[conjugate acid]/[base]}
4.6= pKb + log {0.8/0.5}
pKb = 4.396


use:
pKb = -log Kb
4.396= -log Kb
Kb = 4.018*10^-5
Answer: 4.02*10^-5